OpenStudy (anonymous):

A doctor prescribes an ointment thats is 2% hydrocortisone. A pharmacist has 1% and 5% concentrations in stock. How many ounces of each should the pharmacist use to make a 1 ounce tube?

OpenStudy (anonymous):

lets call the number of ounces of 5% solution \(x\) so the ounces of 2% solution must be \(1-x\) (since the total is one once)

OpenStudy (anonymous):

damn that was wrong, it is 1% solution let me start again lets call the number of ounces of 5% solution \(x\)so the ounces of 1% solution must be \(1−x\) (since the total is one once)

OpenStudy (anonymous):

hmm i have .05x + .01y = 1

OpenStudy (anonymous):

then you have \(.05x+.01(1-x)\) as the amount of hydrocortisone, and this has to be 2% of one ounce with is \(0.02\)

OpenStudy (anonymous):

your job is to solve \[.05x+.01(1-x)=.02\] or to make it easier multiply by 100 to clear the annoying decimals and solve \[5x+(1-x)=2\]

OpenStudy (anonymous):

you get \(4x+1=2\) and this is easy to solve

OpenStudy (anonymous):

How many ounces of each?

OpenStudy (anonymous):

isnt that only one answer?

OpenStudy (anonymous):

what did you get for \(x\) ?

OpenStudy (anonymous):

x= 1/4 or .25

OpenStudy (anonymous):

right, and of course we have remember what \(x\) represents, which is why it is good to define your variable at the beginning. first like i wrote was "lets call the number of ounces of 5% solution \(x"\)

OpenStudy (anonymous):

so we can remember what \(x\) is supposed to be, the number of ounces of 5% solution, which turns out to be 0.25 ounces

OpenStudy (anonymous):

then plug that in to the origional equation?

OpenStudy (anonymous):

now that you know the amount of 5% solution, you also know the amount of 1% solution

OpenStudy (anonymous):

since the total is one ounce

OpenStudy (anonymous):

so .25 percent of 5 % solution and .75 of 1 %?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

yw, hope method was clear